# Gaia Data Release 3: spectroscopic binary-star orbital solutions

### v1

@@ 6.3.4 Fit of the circular orbital solution
The module described here is the SB1CRefiner. The best approached period obtained in the PeriodFinder package is refined and fitted along with the linear parameters , and to fit a circular model to data (see equations in Sect. 3.2). When a circular orbit is selected, corresponds to the maximum of velocity by definition. A flag is written when the derived variance-covariance matrix of the circular solution is singular; this is flag nr.16. If the iterative process does not converge after 1000 iterations, a flag is raised (flag nr.21). If the numerical variance-covariance matrix is not semi-definite positive, a flag is raised (flag nr.19).
@@ 6.3.5 Fit of trends
Not all the observed RV curves correspond to a true periodicity nor to a pseudo-one. One widespread alternative possibility is the existence of a mere trend: the data as a function of time can be represented by a polynomial function of (see Sect. 3.3). The last computing package TrendSolution applies polynomials to data to search for acceleration models (trends) rather than periodic orbits. We start by applying a zero degree polynomial (corresponding to a noisy constancy), then we successively increase the order at each step. The lower degree polynomial is thus nested in the new one. The choice between the lower and the higher degree polynomial is done through a classical Snedecor-F test on the ratio of the improvement of the to the value of the for the higher level model, since the polynomials given above are indeed nested models. A Snedecor-F test for nested models is thus applied. The higher degree polynomial is accepted by the Snedecor-F test for a confidence larger than 0.99865 (3- probability). We stop iterating at the 4th order, or when the higher-degree model exhibits more than one extremum over the time baseline. This last criterion has been adopted in order to avoid any possible overlaps with the orbital models with periods between and (see below). Most of the objects entering the trend category are probably long-term periodic curves but with a timescale much longer than the present time span of the mission (). The code is built to accept polynomials of degrees 1, 2, 3, and 4 but in DR3, only degrees 1 and 2 were encountered (at the selected probability threshold). A flag is written when the derived variance-covariance matrix of the trend solution is singular; this is flag nr.17. If the numerical variance-covariance matrix is not semi-definite positive, a flag is raised (flag nr.20). A regrettable mistake occurred in the code and, for the last operational run, the flags raised were not the correct nr.17 and nr.20 but instead flags nr.15 and nr.18, respectively. This problem has been corrected for DR4.
The coefficients of the polynomial are included in the database but the work is performed in an Hipparcos-like formalism (see section 2.3.3, ESA, 1997). The coefficients saved in the database are related to the derivatives of the velocity curve and are expressed with respect to the DR3 reference epoch (J2016.0, equivalent to BJD 2457389.0).
@@
where is the time (in days) relative to the reference epoch. We thus fit to the time-series the polynomial
with potentially the fixed to zero. In the case of the first degree polynomial ( = 0), we adopt the following rigorous formulae
The and parameters are not correlated if . This is approximated by subtracting to the BJD dates the reference epoch (J2016.0) as a proxy of the middle of the current mission. Practically, not all actual time-series are centred on this value since some transits might be missing (e.g. more numerous before the mid-epoch or alternatively after it). This is in any case still an approximation since it is implicitly assumed that the distribution of the sample times are homogeneously apportioned. This is certainly not true, particularly if one includes EPSL data.
@@
where is the total duration of the mission treated in the framework of DR3. We thus adopt the approximate formulae
@@ 6.3.6 Selection of the best solution
Actually, the procedure described above has been applied when orbital solutions presented a period lower than ; it concerns the regime where the orbital motion is expected, when selected, to be convincingly robust. For longer periods, the folding with the period does not permit to test the perfect overlap homogeneously at all phases. We found it preferable to adopt an alternative procedure. If the orbital period is comprised between and , we decided to enforce the choice of the orbital solution. For periods larger than , i.e. outside the zone of possible convincing orbital motion, we decided to enforce the choice of the TrendSB1 solution. This last point explains why we decided to explore the periodogram in the range between and .
@@ 6.5 Flagging
Beyond the determination of the various parameters, the code also produces various flags that allow the user to have a good understanding of the results and to select some solutions that could be better or more secure than others. The flags generated by our NSS-SB1 processing are incorporated into a more global set of flags issued from the global NSS processing. We will restrict the discussion here to the set of flags directly related to the SB1 processing chain, with a few information linked to flags issued from the SB2 processing that is fully described in Damerdji et al. (2024). All the NSS flags are included in a unique long variable, whose every bit acts as an individual logical flag. All the flags involved in the NSS spectroscopic processing are enumerated in Appendix I. As an example, the SB1Refiner processes the time-series of an object and produces a singular matrix. The flag nr.15 is thus raised, giving to the global flag an additive value of 32768. The object is then processed by the SB1CRefiner and, let say, it concluded to a singular matrix associated to the circular orbit. The flag nr.16 is thus raised (associated to a value of 65536). Consequently, the object is associated to both flags nr.15 and nr.16 raised, giving together a value of 32768+65536=98304. The object could appear in the catalogue as a TrendSB1 with the two above-mentioned flags raised. Although the majority of the flags are important for the decision tree of the algorithm, some of them are purely indicative. This is the case of the flag nr.24 that draws the attention on the fact that the attributed eccentric orbit appears, in the eccentricity versus log diagram, outside the zone where the circularisation of the orbit could have partly taken place. The borderline used is inspired from the one defined by Halbwachs et al. (2005) but modified by us333Using the formalism of Halbwachs et al. (2005), we estimated a value of on the basis of a fairly crude analysis of the full SB9 catalogue (Pourbaix et al. 2004). This estimate has been deduced without proper treatment of the dispersion of due to noise and is considered as a conservative border. The purpose is to flag extremely pathological cases to trigger the cautiousness of the general user.. Flag nr.25 just indicates that the selected adopted period was compatible with one period suggested by the photometry. Flag nr.22 informs that the Lucy test has been applied. Flags nr.11 and 12 concern SB2 objects that turned out to be impossible to treat as SB2 either because the amount of RVs was too small for the secondary (but not for the primary) or because the period found by the SB1 chain was not coherent with the one of the SB2 chain (see Damerdji et al. 2024). Flag nr.26 is raised by the SB2 processing to indicate that the pair of RVs of the SB2 measurements were not anticorrelated (more details are given in Damerdji et al. 2024).
7 The post-processing
@@ 7.1 Internal filtering
In order to maintain the quality level of the various solutions, we defined an internal filtering of these solutions that will allow keeping the very best and most secure ones. The aim is also to spot out and reject solutions that are problematic. Most of the filtering is done via the definition of thresholds. Very often, the selection is made in a statistical way and by using probability thresholds based on the distribution of some statistical variables. As always, the selection near the borders is never perfect. We detail here the internal filtering for SB1 solutions.
@@
In the course of the NSS-SB1 processing, various flags are raised, tracking the details of the decision tree and of the problems encountered. As explained in Sect. 6.5, some combinations of flags imply that the solution is not acceptable or is presumably inaccurate or too hazardous. The corresponding selection was of course considered. Since the internal filtering is supposed to protect the results from spurious solutions and that, in the present purely spectroscopic SB-subcatalogue, no confirmation comes from other channels, we decided to be highly cautious for this step. For the SB1 solutions, we discarded objects having at least one of the three flags nr.14, nr.15, nr.18 raised. As a consequence, this suppresses all the SB1 solutions issued by the SB1CRefiner.
@@
When the supposedly periodic RV data are folded in phase, it happens that a zone in the phase diagram is not populated by any data points. If this region is too wide, the corresponding solution could be pathological with respect to the sampling. Therefore, we choose to discard the solution with a too large where this value represents the largest interval in phase with no data points. We reject solutions leading to . The value of the gap for remaining solutions is not propagated to the SB-subcatalogue but will be accessible in DR4 with the release of the epoch RVs. The threshold value itself has been optimised on the basis of the various validation runs (see Sect. 9 and in particular Sect. 9.2). Due to a small bug in our software, a small part of the objects presenting have not been properly rejected in DR3.
The mere existence of the periodicity in the data is a strict test of the validity of the orbital solution. It is reasonable to consider values for the probability conf_spectro_period from 0.95 to 0.995 or even stricter. During the validation, it appeared a strong population with short-period orbital solutions sometimes associated to relatively large eccentricities. This is somewhat in contradiction with the expected circularisation phenomenon (or these are not properly speaking orbital solutions). Therefore, we decided again to be conservative and to put a threshold that is period dependent. We hope to improve this procedure for DR4. For periods above 10 days, we selected a lower threshold of 0.95 on conf_spectro_period (significance level 0.05). For periods less than one day, we adopted 0.995 as the lower limit. In-between, the lower threshold on conf_spectro_period is given by . The solutions associated to a value less than 0.95 are filtered out. The solutions associated to a value above 0.995 are kept. Solutions in-between are kept but a flag (nr.13, i.e. 8192, see Appendix I) is raised if they are below the defined threshold.
Another natural statistic is the spectroscopic significance defined as . This ratio should not be confused with the probabilistic quantity canonically named significance level. A high significance is necessary to adopt a solution. However, one should not forget that is itself an estimator and that statistical fluctuations could give a value of that is exceedingly small, boosting the significance to very large values. The opposite case is also true. We decided to keep the significance as a possible sorting criterion for the future data applications. Therefore, we only rejected all the solutions with a significance lower than 5. This threshold has been fixed for the internal filtering, but it could be advisable to the general user to select a value of the order of 20 to 50 depending on the details of the scientific application. An example of this way of fine-tuning the selection is illustrated in Gaia Collaboration et al. (2023a).
@@
An internal filtering was similarly defined for the TrendSB1 solutions. This last filtering was on , on the bad flags and on the equivalent of the above-mentioned gap in phase. Concerning bad flags for TrendSB1 solutions, we should have discarded objects with either flag nr.17 or flag nr.20 raised. However, these flags were never activated. We decided to filter out on the flags nr.15 and nr.18, although there could be some unfortunate interactions with the SB1 channel. We rejected solutions presenting a gap (in time) in the time-series larger than 0.3.
@@ 7.2 Combined solutions
After the spectroscopic processing described here above, the solutions are delivered to the combiner. The combiner and combinations are described elsewhere (see section 7.7 by Gavras & Arenou in Pourbaix et al. 2022) and are not extensively addressed here. This combiner explores opportunities to unveil solutions common to other channels. In the framework of DR3, two possibilities were considered.
Either the NSS combines objects from the spectroscopic channel with those issued from the photometric channel and present in the class EclipsingBinary, i.e. presenting eclipses in photometry (see section 7.6 by Siopis & Sadowski in Pourbaix et al. 2022; Siopis et al. 2024; Gaia Collaboration et al. 2023a). The main criterion for the combination is the existence of a common value for the period. If the tentative combination is successful, a global fit over the photometric and the spectroscopic data is performed. In such a case, the new (improved) parameters supersede the old one and are introduced in the catalogue in the new class EclipsingSpectro (see section 7.7 by Gavras & Arenou in Pourbaix et al. 2022). Most of the contributors to this class are issued from our set of SB1 solutions, although some other types of solutions could also be considered (SB1C, SB2, SB2C).
Alternatively, NSS can try to find objects for which the results of the astrometric channel (Halbwachs et al. 2023; Gaia Collaboration et al. 2023a) can be combined with objects having a solution issued from the spectroscopic pipeline. The relevant combined solutions and the related objects are then redirected towards the new class labelled AstroSpectroSB1 (see section 7.7 by Gavras & Arenou in Pourbaix et al. 2022). New fits are here also performed and the old values of the parameters are also superseded. We recall that, in this case, the fit in DR3 is done in the framework of the formalism of Thieles-Innes (Halbwachs et al. 2023; Gaia Collaboration et al. 2023a). This formalism implies that in DR3 AstroSpectroSB1, the combined solutions include two different values (one astrometric and one spectroscopic) for the argument of periastron. The combination with astrometry accepts several types of possible spectroscopic input (SB1, SB1C, TrendSB1, StochasticSB1).
The objects corresponding to a successfully combined spectroscopic solution are removed from the spectroscopic output. Even after the combination process, it might be that an object appears in two classes. In the framework of DR3, these are the sole allowed/activated possibilities. The criteria to consider a combination opportunity are described in details in section 7.7 of Gaia archive documentation444https://gea.esac.esa.int/archive/documentation/GDR3/ (Pourbaix et al. 2022).
@@
In order to avoid losing interesting objects and in a purely prototypical context, the combination was performed before the internal filtering. It could have been more secure to perform it after the internal filtering. Also, the combination of the astrometric solution with the StochasticSB1 one should certainly be a matter of concern. The StochasticSB1 solutions by essence present no actual periodicity and are not expected to be further combined (in principle). On the opposite, solutions that are purely spectroscopic and are ultimately populating the classes SB1, SB1C and TrendSB1 were necessitating increased caution and were submitted to the internal filtering.

### v2

@@ 6.3.4 Fit of the circular orbital solution
The module described here is the SB1CRefiner. The best approached period obtained in the PeriodFinder package is refined and fitted along with the linear parameters , and to fit a circular model to data (see the equations in Sect. 3.2). When a circular orbit is selected, corresponds to the maximum of velocity by definition. A flag is written when the derived variance-covariance matrix of the circular solution is singular; this is flag nr.16. If the iterative process does not converge after 1000 iterations, a flag is raised (flag nr.21). If the numerical variance-covariance matrix is not semi-definite positive, a flag is raised (flag nr.19).
@@ 6.3.5 Fit of trends
Not all the observed RV curves correspond to a true periodicity nor to a pseudo-one. One widespread alternative possibility is the existence of a mere trend: the data as a function of time can be represented by a polynomial function of (see Sect. 3.3). The last computing package TrendSolution applies polynomials to data to search for acceleration models (trends) rather than periodic orbits. We start by applying a zero degree polynomial (corresponding to a noisy constancy), then we successively increase the order at each step. The lower degree polynomial is thus nested in the new one. The choice between the lower and the higher degree polynomial is done through a classical Snedecor-F test on the ratio of the improvement of the to the value of the for the higher level model, since the polynomials given above are indeed nested models. A Snedecor-F test for nested models is thus applied. The higher degree polynomial is accepted by the Snedecor-F test for a confidence larger than 0.99865 (3- probability). We stop iterating at the 4th order, or when the higher-degree model exhibits more than one extremum over the time baseline. This last criterion has been adopted to avoid any possible overlaps with the orbital models with periods between and (see below). Most of the objects entering the trend category are probably long-term periodic curves but with a timescale much longer than the present time span of the mission (). The code is built to accept polynomials of degrees 1, 2, 3, and 4 but in DR3, only degrees 1 and 2 were encountered (at the selected probability threshold). A flag is written when the derived variance-covariance matrix of the trend solution is singular; this is flag nr.17. If the numerical variance-covariance matrix is not semi-definite positive, a flag is raised (flag nr.20). A regrettable mistake occurred in the code and, for the last operational run, the flags raised were not the correct nr.17 and nr.20 but instead flags nr.15 and nr.18, respectively. This problem has been corrected for DR4.
The coefficients of the polynomial are included in the database but the work is performed in an Hipparcos-like formalism (see Sect. 2.3.3, ESA, 1997). The coefficients saved in the database are related to the derivatives of the velocity curve and are expressed with respect to the DR3 reference epoch (J2016.0, equivalent to BJD 2 457 389.0).
@@
where is the time (in days) relative to the reference epoch. We thus fit to the time series the polynomial,
with the value potentially fixed to zero. In the case of the first degree polynomial ( = 0), we adopt the following rigorous formulae,
The and parameters are not correlated if . This is approximated by subtracting to the BJD dates the reference epoch (J2016.0) as a proxy of the middle of the current mission. Practically, not all actual time series are centred on this value since some transits might be missing (e.g. more numerous before the mid-epoch or alternatively after it). This is in any case still an approximation since it is implicitly assumed that the distribution of the sample times are homogeneously apportioned. This is certainly not true, particularly if we include the EPSL data.
@@
where is the total duration of the mission treated in the framework of DR3. We thus adopt the approximate formulae,
@@ 6.3.6 Selection of the best solution
Actually, the procedure described above has been applied when orbital solutions presented a period lower than ; it concerns the regime where the orbital motion is expected, when selected, to be convincingly robust. For longer periods, the folding with the period does not permit us to test the perfect overlap homogeneously at all phases. We found it preferable to adopt an alternative procedure. If the orbital period is comprised between and , we decided to enforce the choice of the orbital solution. For periods larger than , i.e. outside the zone of possible convincing orbital motion, we decided to enforce the choice of the TrendSB1 solution. This last point explains why we decided to explore the periodogram in the range between and .
@@ 6.5 Flagging
Beyond the determination of the various parameters, the code also produces various flags that allow the user to have a good understanding of the results and to select some solutions that could be better or more secure than others. The flags generated by our NSS-SB1 processing are incorporated into a more global set of flags issued from the global NSS processing. We will restrict the discussion here to the set of flags directly related to the SB1 processing chain, with a few information linked to flags issued from the SB2 processing that is fully described in Damerdji et al. (2024). All the NSS flags are included in a unique long variable, whose every bit acts as an individual logical flag. All the flags involved in the NSS spectroscopic processing are enumerated in Appendix I. As an example, the SB1Refiner processes the time series of an object and produces a singular matrix. The flag nr.15 is thus raised, giving to the global flag an additive value of 32 768. The object is then processed by the SB1CRefiner and, let say, it concluded to a singular matrix associated to the circular orbit. The flag nr.16 is thus raised (associated to a value of 65 536). Consequently, the object is associated to both flags nr.15 and nr.16 raised, giving together a value of 32 768 + 65 536 = 98 304. The object could appear in the catalogue as a TrendSB1 with the two above-mentioned flags raised. Although the majority of the flags are important for the decision tree of the algorithm, some of them are purely indicative. This is the case of the flag nr.24 that draws the attention to the fact that the attributed eccentric orbit appears, in the eccentricity versus log diagram, outside the zone where the circularisation of the orbit could have partly taken place. The borderline used is inspired from the one defined by Halbwachs et al. (2005) but modified by us333Using the formalism of Halbwachs et al. (2005), we estimated a value of on the basis of a fairly crude analysis of the full SB9 catalogue (Pourbaix et al. 2004). This estimate has been deduced without proper treatment of the dispersion of due to noise and is considered as a conservative border. The purpose is to flag extremely pathological cases to trigger the cautiousness of the general user.. Flag nr.25 just indicates that the selected adopted period was compatible with one period suggested by the photometry. Flag nr.22 informs that the Lucy test has been applied. Flags nr.11 and 12 concern SB2 objects that turned out to be impossible to treat as SB2 either because the amount of RVs was too small for the secondary (but not for the primary) or because the period found by the SB1 chain was not coherent with the one of the SB2 chain (see Damerdji et al. 2024). Flag nr.26 is raised by the SB2 processing to indicate that the pair of RVs of the SB2 measurements were not anticorrelated (more details are given in Damerdji et al. 2024).
7 Post-processing
@@ 7.1 Internal filtering
In order to maintain the quality level of the various solutions, we defined an internal filtering of these solutions that would allow us to retain the very best and most secure ones. The aim is also to spot out and reject solutions that are problematic. Most of the filtering is done via the definition of thresholds. Very often, the selection is made in a statistical way and by using probability thresholds based on the distribution of some statistical variables. As always, the selection near the borders is never perfect. We detail here the internal filtering for SB1 solutions.
@@
In the course of the NSS-SB1 processing, various flags are raised, tracking the details of the decision tree and of the problems encountered. As explained in Sect. 6.5, some combinations of flags imply that the solution is not acceptable or is presumably inaccurate or too hazardous. The corresponding selection was of course considered. Since the internal filtering is supposed to protect the results from spurious solutions and that, in the present purely spectroscopic SB subcatalogue, no confirmation comes from other channels, we decided to be highly cautious for this step. For the SB1 solutions, we discarded objects having at least one of the three flags nr.14, nr.15, nr.18 raised. As a consequence, this suppresses all the SB1 solutions issued by the SB1CRefiner.
@@
When the supposedly periodic RV data are folded in phase, it happens that a zone in the phase diagram is not populated by any data points. If this region is too wide, the corresponding solution could be pathological with respect to the sampling. Therefore, we choose to discard the solution with a too large where this value represents the largest interval in phase with no data points. We reject solutions leading to . The value of the gap for remaining solutions is not propagated to the SB subcatalogue but will be accessible in DR4 with the release of the epoch RVs. The threshold value itself has been optimised on the basis of the various validation runs (see Sect. 9 and in particular Sect. 9.2). Due to a small bug in our software, a small part of the objects presenting have not been properly rejected in DR3.
The mere existence of the periodicity in the data is a strict test of the validity of the orbital solution. It is reasonable to consider values for the probability conf_spectro_period from 0.95 to 0.995 or even stricter. During the validation, it appeared a strong population with short-period orbital solutions sometimes associated to relatively large eccentricities. This is somewhat in contradiction with the expected circularisation phenomenon (or these are not properly speaking orbital solutions). Therefore, we decided again to be conservative and to put a threshold that is period dependent. We hope to improve this procedure for DR4. For periods above 10 days, we selected a lower threshold of 0.95 on conf_spectro_period (significance level 0.05). For periods less than one day, we adopted 0.995 as the lower limit. In-between, the lower threshold on conf_spectro_period is given by . The solutions associated to a value less than 0.95 are filtered out. The solutions associated to a value above 0.995 are kept. Solutions in-between are kept but a flag (nr.13, i.e. 8192; see Appendix I) is raised if they are below the defined threshold.
Another natural statistic is the spectroscopic significance defined as . This ratio should not be confused with the probabilistic quantity canonically named significance level. A high significance is necessary to adopt a solution. However, we should not forget that is itself an estimator and that statistical fluctuations could give a value of that is exceedingly small, boosting the significance to very large values. The opposite case is also true. We decided to keep the significance as a possible sorting criterion for the future data applications. Therefore, we only rejected all the solutions with a significance lower than 5. This threshold has been fixed for the internal filtering, but it could be advisable to the general user to select a value of the order of 20 to 50 depending on the details of the scientific application. An example of this way of fine-tuning the selection is illustrated in Gaia Collaboration et al. (2023a).
@@
An internal filtering was similarly defined for the TrendSB1 solutions. This last filtering was on , on the bad flags and on the equivalent of the above-mentioned gap in phase. Concerning bad flags for TrendSB1 solutions, we should have discarded objects with either flag nr.17 or flag nr.20 raised. However, these flags were never activated. We decided to filter out on the flags nr.15 and nr.18, although there could be some unfortunate interactions with the SB1 channel. We rejected solutions presenting a gap (in time) in the time series larger than 0.3.
@@ 7.2 Combined solutions
After the spectroscopic processing described here above, the solutions are delivered to the combiner. The combiner and combinations are described elsewhere (see Sect. 7.7 by Gavras & Arenou in Pourbaix et al. 2022) and are not extensively addressed here. This combiner explores opportunities to unveil solutions common to other channels. In the framework of DR3, two possibilities were considered.
Either the NSS combines objects from the spectroscopic channel with those issued from the photometric channel and present in the class EclipsingBinary, i.e. presenting eclipses in photometry (see Sect. 7.6 by Siopis & Sadowski in Pourbaix et al. 2022; Siopis et al. 2024; Gaia Collaboration et al. 2023a). The main criterion for the combination is the existence of a common value for the period. If the tentative combination is successful, a global fit over the photometric and the spectroscopic data is performed. In such a case, the new (improved) parameters supersede the old one and are introduced in the catalogue in the new class EclipsingSpectro (see Sect. 7.7 by Gavras & Arenou in Pourbaix et al. 2022). Most of the contributors to this class are issued from our set of SB1 solutions, although some other types of solutions could also be considered (SB1C, SB2, SB2C).
Alternatively, NSS can try to find objects for which the results of the astrometric channel (Halbwachs et al. 2023; Gaia Collaboration et al. 2023a) can be combined with objects having a solution issued from the spectroscopic pipeline. The relevant combined solutions and the related objects are then redirected towards the new class labelled AstroSpectroSB1 (see Sect. 7.7 by Gavras & Arenou in Pourbaix et al. 2022). New fits are here also performed and the old values of the parameters are also superseded. We recall that in this case, the fit in DR3 is done in the framework of the formalism of Thieles-Innes (Halbwachs et al. 2023; Gaia Collaboration et al. 2023a). This formalism implies that in DR3 AstroSpectroSB1, the combined solutions include two different values (one astrometric and one spectroscopic) for the argument of periastron. The combination with astrometry accepts several types of possible spectroscopic input (SB1, SB1C, TrendSB1, StochasticSB1).
The objects corresponding to a successfully combined spectroscopic solution are removed from the spectroscopic output. Even after the combination process, it might be that an object appears in two classes. In the framework of DR3, these are the sole allowed and/or activated possibilities. The criteria to consider a combination opportunity are described in details in Sect. 7.7 of Gaia archive documentation444https://gea.esac.esa.int/archive/documentation/GDR3/ (Pourbaix et al. 2022).
@@
To avoid losing any interesting objects (and in a purely prototypical context), the combination was performed before the internal filtering. It could have been more secure to perform it after the internal filtering. Also, the combination of the astrometric solution with the StochasticSB1 one should certainly be a matter of concern. The StochasticSB1 solutions by essence present no actual periodicity and are not expected to be further combined (in principle). On the opposite, solutions that are purely spectroscopic and are ultimately populating the classes SB1, SB1C and TrendSB1 were necessitating increased caution and were submitted to the internal filtering.